Angela has $90 to spend on movies and books. The price of movies (M) is $10 and the price of books (B) is $15. Which of the following one or more equations describe Angela's budget constraint?
10M + 15B = 90
9/M = 1+ (3B/2M)
2M = 18 - 3B
B = 6 + (2/3)M
Answer #1

Angela's budget constraint is :

10M + 15B = 90

Where $90 is the money income, $10 is the price of movies & $15 is the price of books. Also the budget constraint can be written as :

(5×2)M + (5×3)B = 90

Taking 5 common

5(2M + 3B) = 90

2M + 3B = 18

2M = 18 - 3B